# The Peripatetic Hater: Predicting Movement Among Hate Subreddits

### v1

@@ Discussion
This study provides a comprehensive view of the user dynamics among multiple types of hate communities on Reddit. These results alone are remarkable as they suggest that hate manifested in the current historical moment among English-speaking users of this platform can be objectively categorized into three dimensions. But why three? The dimension we term “Racist” is associated with racism, xenophobia, and antisemitism, which are types of hate that stem from an essentialist point of view, or the false notion that there are inherent biological differences among people of different races or national origins (Roth, van Stee, and Regla-Vargas 2023). The second dimension, in contrast, is associated with homophobia, transphobia, and, to a lesser degree, misogyny, i.e., hate associated with perceived choice (the false belief being that being LGBTQ is not an immutable characteristic of the individual, but rather is under their control). Finally, the third category is misogyny and, to a lesser extent, fat shaming (as we observe fat shaming is often directed at women (Bound Alberti 2021), and is associated with the objectification of women), which is hate associated with gender. These categories are intrinsically disparate types of hate even though the identities attacked are diverse within each category. How this generalizes to other social media, and how it relates to psychological perceptions of hate, constitute rich avenues for future research.
@@
We have also shown that users who initially participate in one hate community are prone to participate in additional categories of hate communities and that participating in additional categories of hate is associated with greater activity in hate communities and a broader hate group lexicon. We can, however, better predict which users will join additional communities compared to baselines. This may be in part because peripatetic users employ more of the language from within the community that they will eventually participate in. Below we discuss the possible implications of these results concerning how hate communities can influence each other online.
RQ1: users likely to participate in multiple communities of hate. Figure 3 shows that users who originate in one type of hate community are more likely to participate in an alternate type of hate community than are similar users who did not participate in the original hate community type. Moreover, the likelihood of joining a given alternate type of hate community does not differ depending on the original type (Figure 4A). Taken together, these results imply that each type of hate community serves as a “gateway” to other hate communities. In other words, the presence of each type of community could further amplify the growth of the others, as their user bases appeal to each other. Future work can trace the reasons why users from each category of hate are more likely to join other hate subreddits – for example, the impact of recruitment on peripatetic behavior can be studied (Russo, Ribeiro, and West 2023).
@@
RQ2: characteristics distinguishing peripatetic and non-peripatetic users. Given the results displayed in Figure 4B, it is tempting to conclude that joining additional hate groups leads peripatetic users to import hate terms and ideas when posting in the subreddit in which they first became active. We explored this possibility, but could not find conclusive evidence of the causal direction. Two non-mutually-exclusive explanations therefore seem plausible. First, it is possible that users who become peripatetic were interested from the outset in the topics of the additional subreddits that they eventually joined, i.e., differences in language use between users who do and do not subsequently become peripatetic reflect differences in the breadth of their antipathies at time zero. Second, users who eventually post in other subreddits may have been consuming content from those subreddits long before they first contributed to them, hence said content may have influenced their posts in the first subreddit which they joined. Regardless of which of these various possibilities applies, given that a) social influence likely occurs, and b) consuming content from, and contributing content to, multiple communities is facilitated by their presence on a single platform, the harm posed by hosting multiple types of hate communities on one platform may be greater than the sum of the harm posed by each community. In this scenario, hate subreddits serve as recruitment venues for each other, both enhancing the growth of their respective user bases and leading to a convergence of hate beliefs over time.
RQ3: Joining additional hate communities may further radicalize users. Figure 4C–D demonstrates the change in user behavior as they join additional hate subreddits. Namely, we see an increase in activity and a broader lexicon of these users, especially if they join subreddits of a different category. This may imply that users become more active in, and internalize language of, these various dimensions of hate. These features could be a first sign that users may radicalize more quickly by joining additional hate groups (Schmitz et al. 2024).
@@
RQ4: Predicting peripatetic movement. Table 1 illustrates our model’s prediction performance for the types of hate subreddits users that will eventually join. The relatively poor performance of the model suggests that stochastic or unobservable factors strongly influence who becomes a peripatetic hater. Nonetheless, for each category, performance is above the baseline, indicating that peripateticism is not a purely spurious decision. Users are either partly motivated to join based on their past multi-hate category opinions represented in writing, or their writing indicates that they are more vulnerable to being influenced to join other types of hate. This technology is the first step towards understanding users’ motivations for joining multiple hate subreddits and sheds light on where interventions designed to reduce the harm of hate subreddits might most productively focus.
@@
Predictive modeling. The model we use is limited in its efficacy to predict where users will join. As mentioned previously, we cannot rule out the possibility that joining additional subreddits is a relatively stochastic decision or is affected by unobservable factors. Future work will need to explore motivations for joining additional hate subreddits in greater detail. Should a more predictive model become possible, however, this technology could help platforms develop proactive content moderation strategies by reducing problematic content on social media before it has been created (Habib et al. 2021). Moreover, while we employ techniques such as topic modeling to understand the textual features that are associated with users’ contributions to different hate groups, these techniques do not provide direct insight into why our model produced the predictions that it did. Existing methods and further advancements in the field of explainable natural language processing can help us understand how our predictive model learns to classify users as peripatetic or non-peripatetic (Zini and Awad 2022).
How peripatetic users contribute to subreddit bans. Previous research shows that the proportion of members of a subreddit who previously posted in banned or quarantined subreddits is strongly predictive of the subreddit being banned or quarantined in the future (Habib et al. 2021). In order to uncover the mechanisms by which peripatetic users contribute to given subreddits being banned, future work could more deeply investigate the activity of peripatetic users temporally in relation to the ban dates of each subreddit. Are these users simply interested in subreddits that are already likely to be banned, or does their activity within the subreddit directly or indirectly contribute to subreddit bans? Within each subreddit, one could identify large influxes of users who have already posted in other hate subreddits. Network analyses could be used to investigate the diffusion of ideas from peripatetic to non-peripatetic users, and dynamic topic modeling could be used to understand the shift in topics discussed by the core user bases of the subreddits over time (Wang, Blei, and Heckerman 2012). Such work could ultimately help to determine whether peripatetic users introduce problematic narratives to subreddits that eventually result in communities being banned.

### v2

@@ Discussion
This study provides a comprehensive view of the user dynamics among multiple types of hate communities on Reddit. These results suggest that hate manifested in the current historical moment among English-speaking users of this platform can be categorized into eight dimensions roughly reflecting each major identity from federal definitions of targeted groups. While these results largely corroborate the federal targets of hate, we acknowledge that there could be yet more identities that we did not capture. How this generalizes to other social media, and how it relates to psychological perceptions of hate, constitute rich avenues for future research.
@@
Importantly, our results indicate that users who initially participate in one hate community are prone to participate in additional categories of hate communities and that participating in additional categories of hate is associated with a broader hate group lexicon. Using AI tools, we are able to predict which users will join additional hate communities compared to baselines. Part of this predictive power may be because, from the outset, users who will become peripatetic employ more of the language associated with the other hate communities that they will eventually join. Below we discuss the possible implications of these results concerning how hate communities can influence each other online.
RQ1: users are likely to participate in multiple communities of hate. Figure 3 shows that users who originate in one type of hate community are more likely to participate in an alternate type of hate community than are similar users who did not participate in the original hate community type. This result implies that each type of hate community serves as a “gateway” to other hate communities. In other words, the presence of each type of community could further amplify the growth of the others, as their user bases appeal to each other. The preferences certain types of hate communities have for each other provides insight into why this may be. For example, racist, xenophobic, antisemitic, and Islamophobic subreddits all have strong user overlap with each other, likely because these types of hate are similar in that they tend to attribute deprecated characteristics to these essences. For example, hate speech directed at Jews and Muslims will often target their perceived race or national origin along with their religious identity.
Future work should explore the reasons why users from each category of hate are more likely to join other hate subreddits – for example, due to recruitment (Russo, Ribeiro, and West 2024). Additional research is also needed to determine whether hate group categorizing could help identify previously unknown active hateful communities or clusters, possibly leading to quicker moderation.
@@
RQ2: characteristics distinguishing peripatetic and non-peripatetic users. Given the results displayed in Figure 4B, it is tempting to conclude that joining additional hate groups leads peripatetic users to import hate terms and ideas when posting in the subreddit in which they first became active. We explored this possibility, but could not find conclusive evidence of the causal direction. Two non-mutually-exclusive explanations therefore seem plausible: 1) users who become peripatetic were interested from the outset in the hate categories of the subreddits that they eventually joined, or 2) users who eventually post in other subreddits may have been consuming content from those subreddits long before they first contributed to them, hence said content may have influenced their posts in the first subreddit that they joined. Regardless of which of these various possibilities applies, given that a) social influence likely occurs, and b) consuming content from, and contributing content to, multiple communities is facilitated by their presence on a single platform, the harm posed by hosting multiple types of hate communities on one platform may be greater than the sum of the harm posed by each community. In this scenario, hate subreddits serve as recruitment venues for each other, both enhancing the growth of their respective user bases and leading to a convergence of hate beliefs over time. Finally, our work on movement complements prior work on understanding the formation of collective identity within these groups (Gaudette et al. 2021). Future work could explore how collective identity affects the likelihood that users become peripatetic.
RQ3: Joining additional hate communities may further radicalize users. Figure 4C shows an increase in peripatetic users’ use of terms associated with the additional categories of subreddits they join. This may imply that users internalize the language of these various dimensions of hate. These features could be a first sign that users may join additional hate groups (Schmitz et al. 2024) and is consistent with, although not proof of, theories that propose social networks act as a net to drive individuals to deeper forms of radicalization (Borum 2011; Maskaliūnaitė 2015; Neo 2016). Furthermore, other theories explore why people join extreme communities in the first place (Mann et al. 2023). Future work is needed to understand whether the motivations for participating in hate communities put forth by these theories play a role in differentiating peripatetic and non-peripatetic hate users.
@@
RQ4: Predicting peripatetic movement. Table 1 illustrates our model’s prediction performance for the types of hate subreddits users that will eventually join. The relatively poor performance of the model suggests that stochastic or unobservable factors strongly influence who becomes a peripatetic hater. Nonetheless, for each category, performance is above the baseline, indicating that peripateticism is not a purely stochastic decision. Users are either partly motivated to join based on their past multi-hate category opinions represented in writing, or their writing indicates that they are more vulnerable to being influenced to join other types of hate. This technology is the first step towards understanding users’ motivations for joining multiple hate subreddits and sheds light on where interventions designed to reduce the harm of hate subreddits might most productively focus.
In sum, our exploration of the dynamics of users in Reddit hate communities reveals the existence of both hazards and opportunities. On the one hand, these dynamics suggest that, over time, online hate groups might converge or coalesce, potentially both accelerating their growth and increasing the probability of tangible offline harms. On the other hand, understanding the movement of users among hate communities and how peripatetic users introduce language that cross-fertilizes hate groups may aid in developing avenues for both identification and intervention aimed at inhibiting the growth and coalescence of hate communities, thereby curtailing the spread of hateful beliefs and reducing the risk of offline harms.
@@
While our study effectively categorizes hate subreddits and demonstrates processes of content diffusion across these categories, there are still notable limitations to our work. First, our hate detection methodology could be improved, especially for transphobia. Our results may also vary with alternative models we did not test, and these results may not generalize to non-English hate subreddits. Our model also lacks exhaustive analysis of hyperparameter tuning. While we found that the variations explored so far do not substantially change results, a more exhaustive analysis could improve the model further. Finally, while we filter out likely bots from our dataset, we do not distinguish between users who post in support of a group’s ideology and users who employ counterspeech (Mathew et al. 2019).
Future work can expand on our results by exploring motivations for joining additional hate subreddits in greater detail. For example, a stronger and more explainable predictive model, which explains why the model made the predictions it did, could help platforms develop proactive content moderation strategies by reducing problematic content on social media before it has been disseminated (Habib et al. 2021). Additionally, as prior research shows that the proportion of users in a given subreddit who previously posted in banned subreddits is predictive of whether that subreddit will be banned in the future (Habib et al. 2021), future work can more closely investigate how peripatetic users may contribute to subreddit bans.
Acknowledgements
The authors would like to thank Hillel Cogan, Emily Haddad, Derrick Liu, Daniel Penn, Rafael Arellano, and Dan Faltesek for their valuable work annotating hate speech. Our work is supported by NSF award #2331722.
Appendix
Cluster
Num.
subreddits
Num. users
Num.
comments
Num.
submissions
Avg. comments
per user
Avg. submissions
per user
General Hate 58 1,623,808 2,716,821 434,951 1.67 0.27
Misogynistic 32 818,835 1,655,450 229,105 2.02 0.28
Anti-LGBTQ 16 54,341 149,444 17,290 2.75 0.32
Racist 15 55,554 140,491 15,450 2.53 0.28
Xenophobic 35 1,577,444 3,445,695 477,587 2.18 0.30
Islamophobic 7 113,770 242,277 29,087 2.13 0.26
Antisemitic 4 10,174 19,433 4,957 1.91 0.49
Ableist 1 2,770 4,183 1,519 1.51 0.55
Table 2: Subreddit cluster statistics
F1
Category Demux BERT
Racism 0.68 $\pm$ 0.005 0.67 $\pm$ 0.001
Misogyny 0.72 $\pm$ 0.001 0.71 $\pm$ 0.008
Xenophobia 0.67 $\pm$ 0.017 0.65 $\pm$ 0.009
Transphobia 0.31 $\pm$ 0.055 0.23 $\pm$ 0.068
Antisemitism 0.62 $\pm$ 0.028 0.54 $\pm$ 0.033
Islamophobia 0.63 $\pm$ 0.022 0.61 $\pm$ 0.025
Homophobia 0.71 $\pm$ 0.005 0.68 $\pm$ 0.006
Ableism 0.69 $\pm$ 0.010 0.66 $\pm$ 0.028
Table 3: Targeted identity prediction performance on the measuring hate speech corpus.
$R^{2}$
Category Demux (maximize $R^{2}$) Demux (maximize F1) GPT-4 (batches) GPT-4 (individual posts)
Racism 0.91 0.83 0.89 0.60
Misogyny 0.87 0.50 0.55 0.29
Xenophobia 0.91 0.35 0.24 0.34
Transphobia 0.52 0.09 0.20 0.01
Antisemitism 0.79 0.79 0.87 0.00
Islamophobia 0.65 0.32 0.37 0.00
Homophobia 0.82 0.78 0.64 0.27
Ableism 0.72 0.29 0.48 0.24
Table 4: Performance of hate speech detection models on annotated subreddit data
Negative Disrespectful Insult Attack Hate speech
Negative - 0.94 0.95 0.97 0.83
Disrespectful - 0.99 0.93 0.87
Insult - 0.95 0.88
Attack - 0.84
Hate speech -
Table 5: Correlation matrix of different hate speech detection methods
Additional robustness checks
GPT-4 Prompting
In addition to using Demux to detect targeted identities in hate subreddits, we explored the possibility of using GPT-4, as it has been shown to perform well on the task of hate speech detection (Kumarage, Bhattacharjee, and Garland 2024). For each type of hate speech, we prompt GPT-4 to count the number of posts containing that type of hate speech in batches of 10 with the following string:
[LIST OF POSTS]
Additionally, we test the performance of a prompt designed to detect individual posts:
Below is a Reddit post that may or may not be attacking someone because of their [BIAS CATEGORY]. Please respond with “Yes” or “No.” Is this post very clearly [TYPE OF HATE SPEECH]? Do not include anything else in your response:
[POST]
Table 4 shows the performance of both these approaches on the set of hate subreddits we manually annotated for hate speech. Across the board, GPT-4 performs better when predicting the number of posts containing hate speech rather than predicting whether individual posts contain hate speech. While GPT-4 exhibits comparable performance to Demux, we opt to use Demux as it is less expensive.
Subreddit clusters using thresholds that optimize $R^{2}$.
As mentioned earlier in the text, we do not know which approach for determining which thresholds to use from our hate speech detection model generalizes better to our broader set of hate communities: using thresholds that maximize $R^{2}$ values on the data we annotated or using the thresholds that maximize the F1 score on the Measuring Hate Speech corpus. For this reason, we repeat our analyses with both methods. When clustering the subreddits using the thresholds that maximize the $R^{2}$ value, we find largely the same categories as the thresholds that maximize the F1 score on the Measuring Hate Speech corpus. The only difference is that the “anti-LGBTQ” cluster is split up into two clusters: a “transphobic” cluster and a “homophobic” cluster. Additionally, we noticed many dark humor subreddits appearing in the transphobic cluster (along with TERF subreddits). These dark humor subreddits are classified as general hate subreddits when using the thresholds that maximize the F1 score. Therefore, as the categorization obtained by the threshold that maximized the F1 score fit better with our ethnographic understanding of the subreddits in our dataset, we chose to report the results obtained from this method in the main text of our paper. The results of all analyses are qualitatively similar between the two methods.
Varying time thresholds for peripatetic activity.
Figure 7: Average effect sizes of Figure 3 when considering different time thresholds for peripatetic users. Effect sizes are calculated as the proportion of joiners that post in an alternate category of hate subreddit within the time threshold divided by the proportion of matched counterparts that post in an alternate category of hate subreddit within the time threshold. Error bars represent standard errors.
Figure 7 shows the average effects of our matched-pairs analysis (Figure 3) when varying the time threshold considered (i.e., what proportion of users post in an alternate hate subreddit within $x$ weeks of posting in their first hate subreddit?). Observing the plot, we can see that the effect size decreases the longer the threshold is. However, when removing the threshold entirely, we find that for all original categories of subreddits, the proportion of joiners who post in an alternate category of hate subreddit remains higher than the proportion of matched counterparts who post in an alternate category. The average effect size when removing the time threshold is 1.4, and all results are statistically significant (p-values $<10^{-10}$).
Figure 8: Transition matrices (Figure 4A) produced when considering different thresholds for considering peripatetic users.
Figure 8 shows the likelihood of a user moving from one category to another within (A) six weeks, (B) six months, and (C) any time at all after initially posting in their first hate subreddit. In all plots, we see similar patterns in which categories of hate subreddits show preferences for each other.
Figure 9: Heatmaps showing relative use of ingroup language by peripatetic vs. non-peripatetic users at different time thresholds. Odds ratios indicate the odds of the use of ingroup language from the alternate category of users who move to the alternate category within the time threshold, relative to the odds of non-peripatetic users using such language.
Figure 9 shows the results of the analysis displayed in Figure 4B at varying time thresholds for considering peripatetic users. The effects are mainly positive for all different time thresholds. Figure 11 shows the results of the same analysis, using posts made at any time before moving to an alternate category of subreddit instead of the first three days (we use the default threshold of six weeks to consider peripatetic users in this case). The results are qualitatively similar to the analysis only considering the first three days of each user’s activity. Similarly, Figure 10 shows the results at different time thresholds for the change in ingroup language use after users become peripatetic. We observe similar results regardless of the time threshold.
Figure 10: Heatmaps showing the change in use of ingroup language by peripatetic users at different time thresholds. Odds ratios indicate the odds of the use of ingroup language from the alternate category after becoming peripatetic, relative to the odds of the use of ingroup language before becoming peripatetic. Users are considered in the sample if they moved to the alternate category within the time threshold displayed on the subplot.
Figure 11: Heatmap showing relative use of ingroup language by peripatetic vs. non-peripatetic users. Unlike Figure 4B, the odds ratios are calculated from all posts made by non-peripatetic users and all posts made by peripatetic users before moving to an alternate category of hate, instead of only posts made in the first three days.
In addition to showing how our results regarding ingroup language use change when considering all posts instead of only the first three days of each user’s activity, we also fit topic models when considering all posts. The top topics are displayed in Figure 12, and the topics with the highest and lowest odds ratios are displayed in Figure 13. In both plots, many of the topics are represented by similar words to the topic plots in the main text. Though there are also plenty of topics that are different, we nonetheless see similar patterns of topics related to alternate categories of hate being positively associated with peripatetic users (for example, the topic in the racist cluster represented by the words “gay” and “homosexuality” being used more frequently by peripatetic users).
Figure 12: Most frequent topics per subreddit category when considering all posts made by users prior to posting in an alternate category of hate subreddit. These data include posts made by peripatetic and non-peripatetic users.
Figure 13: Topics most strongly associated with peripatetic activity when considering all posts made by users prior to posting in an alternate category of hate subreddit. These data include posts made by peripatetic and non-peripatetic users. Topics with an adjusted p-value greater than 0.05 are lighter to show statistical insignificance.